# BianCang: A Traditional Chinese Medicine Large Language Model

### v1

@@ 4.2.2 Subjective Evaluation
We used the BC-Analytical test set for the subjective evaluation of the model. The BC-Analytical dataset, which we constructed using proprietary medical records, includes 50 complex TCM cases. Each case is accompanied by three questions: disease diagnosis and syndrome differentiation analysis, differential diagnosis, and treatment plan formulation. We evaluate the model from three dimensions: professionalism, fluency, and safety, using win rate, tie rate, and loss rate as the evaluation metrics. All evaluations are conducted by hospital physicians. Following prior research [Yang et al., 2024b], we redefined the evaluation criteria for professionalism, fluency and safety. The specific evaluation standards are as follows:
Professionalism: (1) Must have an accurate understanding of the problems and needs from doctors or patients in order to provide relevant responses and advice; (2) must correctly predict the patient’s disease or syndrome and provide supporting rationale when providing diagnostic support; (3) must correctly interpret the medical knowledge involved.
Fluency: (1) Answers must be semantically coherent and free of logical errors or irrelevant information; (2) the style and content of answers must be consistent and free of conflicting information; (3) responses must remain friendly and welcoming; cold or overly brief language is unacceptable.
Safety: (1) Must provide scientifically accurate medical knowledge, especially in cases of disease diagnosis, medication recommendations, etc.; (2) must admit ignorance of unknown knowledge; (3) must ensure patient safety; (4) must refuse to answer information or advice that could cause harm; (5) must comply with medical ethics while respecting patient choice and refuse to answer if violated.
4.3 Results
4.3.1 Objective Evaluation Results
The objective evaluation results in the field of traditional Chinese medicine (TCM) are shown in Table 3. According to the table, BianCang achieved outstanding performance across three TCM tasks: syndrome differentiation, disease diagnosis, and TCM exam. In all cases, it outperformed the baseline models. Below, we summarize our main observations and conclusions from these experimental results:
Most previous large TCM and medical models lacked the analytic capabilities for syndrome differentiation and disease diagnosis in TCM. We identify three main reasons for this limitation: (1) The foundational models of some early models were relatively weak, resulting in insufficient fundamental capabilities; (2) The TCM tasks require systematic knowledge reserves and extensive diagnostic experience, which most previous models lacked due to limited guidance from real-world TCM syndrome differentiation and disease diagnosis experience; (3) Although prior models utilized continuous pre-training and supervised fine-tuning, they did not fully leverage the characteristics of each stage, resulting in an inconsistency between the fine-tuning instructions and the model’s internal knowledge.
Our model possesses strong analytical capabilities for both syndrome differentiation and disease diagnosis. For instance, BianCang-Qwen2.5-7B-Instruct achieved an accuracy of 78.90% on the TCMSD test set in direct inference (DI) mode, reflecting an improvement of approximately 48 percentage points over the foundational model Qwen2.5-7B, and about 47 percentage points over the previously best-performing TCM model, Lingdan-13B. In the chain-of-thought (CoT) reasoning mode, BianCang-Qwen2.5-7B-Instruct achieved further improvement, reaching an accuracy of 82.10% on the TCMSD test set, about 61 percentage points higher than the foundational model Qwen2.5-7B and around 48 percentage points above the previously best-performing TCM model, Huatuo2-7B. Similarly, BianCang-Qwen2.5-7B-Instruct also achieved significant improvements on the TCMSD-BC, TCMDD, and TCMDD-BC test sets.
Moreover, our model achieved optimal results in both TCM and TCM combined with western medicine exams, indicating that it has mastered a more comprehensive and systematic TCM knowledge base. These experimental results verify the effectiveness of our selected foundational model, data formulation, and training strategy within the TCM domain.
The objective evaluation results for the Chinese medical exams are shown in Table 4. As indicated in Table 4, BianCang not only performed exceptionally well in the field of TCM but also demonstrated remarkable performance in modern medicine, surpassing all baseline models. This indicates that our data formulation and training strategy have enhanced the model’s capabilities in both TCM and modern medicine.
@@ 4.3.1 Objective Evaluation Results
Table 5: Performance comparison of one-stage and two-stage training methods. The one-stage approach involves supervised fine-tuning of the vanilla model only. The two-stage approach first performs continuous pretraining on the vanilla model, followed by supervised fine-tuning. The training data and parameter settings for supervised fine-tuning are identical.
4.3.2 Subjective Evaluation Results
For the subjective evaluation, we selected four models with strong overall performance in objective evaluation as baselines to reduce human evaluation costs. The subjective evaluation results of TCM case analysis are shown in Figure 3. The results indicate that our model demonstrates a higher level of professionalism in analyzing real TCM cases, which is largely due to its well-developed TCM knowledge system and authentic diagnostic experience. In terms of fluency, our model can generate more authentic, TCM-styled outputs, surpassing most baseline models. In terms of safety, such as in diagnostic or medication recommendation scenarios, our model consistently outperformed the baseline models. Overall, our model’s responses received greater approval from TCM experts.
@@ 4.4 Effectiveness of Two-stage Training Strategy
We conducted a set of comparative experiments to illustrate the necessity and effectiveness of the two-stage training strategy. The experimental results are shown in Table 5. “Vanilla” refers to the foundational model. “One Stage” represents directly applying supervised fine-tuning (SFT) to the model using an instruction set tailored for TCM downstream tasks. “Two Stage” represents first injecting knowledge through continuous pre-training and then applying supervised fine-tuning on downstream tasks to maximize consistency between the instructions and the model’s internal knowledge during fine-tuning. The SFT data and training parameter settings used in both One Stage and Two Stage are entirely consistent.
@@
The experimental results indicate that our two-stage training strategy is highly effective when the foundational model itself lacks sufficient knowledge. Taking the TCM syndrome differentiation task as an example, we observe that the foundational model performs poorly on this task, highlighting its lack of TCM-specific knowledge in syndrome differentiation. In such cases, directly applying SFT to the base model may result in limited performance gains due to inconsistencies between the instructions and the model’s internal parameter knowledge. In contrast, with the two-stage training strategy, the consistency between the instructions and the model’s internal parameter knowledge is enhanced, leading to a substantial improvement in the model’s performance on the TCM syndrome differentiation task. This improvement is especially significant in chain-of-thought scenarios, where effective reasoning depends on the model’s foundational knowledge.
5 Conclusion and Limitations
@@
In this work, we introduce BianCang, a large language model specifically designed for traditional Chinese medicine (TCM), to address challenges unique to TCM large language models. We constructed a comprehensive pre-training corpus and a diverse set of supervised fine-tuning instructions, employing a two-stage training strategy to train BianCang that first injects domain-specific knowledge and then aligns it through targeted stimulation. Experimental results robustly demonstrate the effectiveness and superiority of BianCang. Despite these achievements, it is important to emphasize that BianCang is primarily an auxiliary research tool, not a substitute for professional TCM consultation. We recognize the limitations of the model. BianCang cannot guarantee accuracy in all its responses. Due to the serious consequences that can arise from inaccurate data in medical domain, we strongly suggest that users exercise caution when dealing with generated information and seek advice from experts.
Acknowledgments
@@
This work was supported by the National Natural Science Foundation of China (Grant No.62376130), Shandong Provincial Natural Science Foundation (Grant No.ZR2022MF243), Program of New Twenty Policies for Universities of Jinan (Grant No.202333008), and Pilot Project for Integrated Innovation of Science, Education, and Industry of Qilu University of Technology (Shandong Academy of Sciences) (Grant No.2024ZDZX08).
We would like to express our heartfelt gratitude to Cong Wang, Kangjun Liu and Jianlei Wang for their invaluable contributions to data processing and experimentation in this work.
Appendix A Evaluation prompt
Table 6: Evaluation prompt
Prompt 1: TCM Syndrome Differentiation Prompt Template
Direct Inference Prompt Template
下面是患者的病历信息。
{主诉}
{病史}
{入院情况}
{中医望闻切诊}
问题:根据患者的上述情况,患者的中医证型为?
答案:
Below is the patient’s medical record information.
{chief complaint}
{medical history}
{admission condition}
{TCM four diagnostic methods}
Question: Based on the patient’s condition above, what is the patient’s Traditional Chinese Medicine (TCM)
syndrome type?
Answer:
Chain of Thought Prompt Template
下面是患者的病历信息。
{主诉}
{病史}
{入院情况}
{中医望闻切诊}
问题:根据患者的上述情况,对该患者进行中医辨证分析,首先给出辨证过程和依据,然后给出患者
的中医证型。
答案:
Below is the patient’s medical record information.
{chief complaint}
{medical history}
{admission condition}
{TCM four diagnostic methods}
Question: Based on the patient’s condition, perform a TCM syndrome differentiation analysis. First, provide the
process and evidence for the differentiation, then identify the patient’s TCM syndrome.
Answer:
Prompt 2: TCM Disease Diagnosis Prompt Template
Direct Inference
下面是患者的病历信息。
{主诉}
{病史}
{入院情况}
{中医望闻切诊}
问题:根据患者的上述情况,患者的中医疾病为?答案中只输出疾病名称,不要输出其它额外文本。
答案:
Below is the patient’s medical record information.
{chief complaint}
{medical history}
{admission condition}
{TCM four diagnostic methods}
Question: Based on the patient’s aforementioned condition, what is the patient’s TCM disease? Only provide
the disease name in your answer, without any additional text.
Answer:
Answer:
Chain of Thought Prompt Template
下面是患者的病历信息。
{主诉}
{病史}
{入院情况}
{中医望闻切诊}
问题:根据患者的上述情况,对该患者进行中医辨病分析,首先给出辨病过程和依据,然后给出患者
的中医疾病。
答案:
Below is the patient’s medical record information.
{chief complaint}
{medical history}
{admission condition}
{TCM four diagnostic methods}
Question: Based on the patient’s condition, perform a TCM disease differentiation analysis. First, provide the
process and evidence for the differentiation, then identify the patient’s TCM disease.
Answer:
Prompt 3: National Medical Licensing Examination in China Prompt Template
以下是关于{问题类型}的一道选择题,不需要做任何分析和解释,直接输出答案选项。
问题:{题目}
选项:A. {选项A} B. {选项B} …
答案:
The following is a multiple-choice question regarding {question type}. No analysis or explanation is required;
simply output the answer option.
Question: {question}
Options: A. {optionA} B. {optionB} …
Answer:
Prompt 4: Chinese Medical Benchmark Prompt Template
以下是中国{考试类型}中{考试科目}考试的一道{问题类型},不需要做任何分析和解释,直接输出
答案选项。
{题目}
A. {选项A}
B. {选项B}
The following is a {question type} from the {exam class} exam in China’s {exam type}. No analysis or explan-
ation is needed, simply output the answer option.
{question}
A. {option A}
B. {option B}
Prompt 5: Case Analysis Prompt Template
以下是一位病人的病例:{case}。
根据病人的病例回答下面的问题。
问题:{question}
答案:
The following is a patient case: {case}.
Based on the patient’s case, answer the following question.
Question: {question}
Answer:

### v2

@@ 4.3 Objective Evaluation Results
Table 5: Performance comparison of one-stage and two-stage training methods. The one-stage approach involves supervised fine-tuning of the vanilla model only. The two-stage approach first performs continual pretraining on the vanilla model, followed by supervised fine-tuning. The training data and parameter settings for supervised fine-tuning are identical.
@@
4.4 Subjective Evaluation Details and Datasets
We used the BC-Analytical dataset for the subjective evaluation of the model. The BC-Analytical dataset, which we constructed using proprietary medical records, includes 50 complex TCM cases. Each case is accompanied by three questions: disease diagnosis and syndrome differentiation analysis, differential diagnosis, and treatment plan formulation. We evaluate the model from three dimensions: professionalism, fluency, and safety, using win rate, tie rate, and loss rate as the evaluation metrics. All evaluations are conducted by two hospital physicians specializing in TCM cardiology. Following prior research [21, 20], we redefined the evaluation criteria for professionalism, fluency and safety. The specific evaluation standards are as follows:
The evaluation criteria for fluency and safety are consistent with previous research [21, 20], and the professionalism evaluation criteria are as follows: (1) Must have an accurate understanding of the problems and needs from doctors or patients in order to provide relevant responses and advice; (2) must correctly predict the patient’s disease or syndrome and provide supporting rationale when providing diagnostic support; (3) must correctly interpret the medical knowledge involved.
4.5 Subjective Evaluation Results
For the subjective evaluation, as shown in Fig.  3, we selected four models with strong overall performance in objective evaluation as baselines to reduce human evaluation costs.
The results indicate that our model demonstrates a higher level of professionalism in analyzing real TCM cases, which is largely due to its well-developed TCM knowledge system and authentic diagnostic experience. In terms of fluency, our model can generate more authentic, TCM-styled outputs, surpassing most baseline models. In terms of safety, such as in diagnostic or medication recommendation scenarios, our model consistently outperformed the baseline models. Overall, our model’s responses received greater approval from TCM experts.
@@ 4.6 Effectiveness of Two-stage Training Strategy
We conducted a set of comparative experiments to illustrate the necessity and effectiveness of the two-stage training strategy. The experimental results are shown in Table 5. “Vanilla” refers to the foundational model, Qwen2.5-7B. “One Stage” represents directly applying supervised fine-tuning (SFT) to the model using an instruction set tailored for TCM downstream tasks. “Two Stage” represents first injecting knowledge through continual pre-training and then applying supervised fine-tuning on downstream tasks to maximize consistency between the instructions and the model’s internal knowledge during fine-tuning. The SFT data and training parameter settings used in both One Stage and Two Stage are consistent.
@@
The experimental results indicate that our two-stage training strategy is highly effective when the foundational model itself lacks sufficient knowledge. Taking the TCM syndrome differentiation task as an example, we observe that the foundational model performs poorly on this task, highlighting its lack of TCM-specific knowledge in syndrome differentiation. In such cases, directly applying SFT to the foundational model may result in limited performance gains due to inconsistencies between the instructions and the model’s internal parameter knowledge. In contrast, with the two-stage training strategy, the consistency between the instructions and the model’s internal parameter knowledge is enhanced, leading to a substantial improvement in the model’s performance on the TCM syndrome differentiation task. This improvement is especially significant in chain-of-thought scenarios, where effective reasoning depends on the model’s foundational knowledge.
5 Conclusion
@@
In this work, we introduce BianCang, a large language model specifically designed for traditional Chinese medicine (TCM), to address challenges unique to TCM large language models. We constructed a comprehensive pre-training corpus and a diverse set of supervised fine-tuning instructions, and employed a two-stage training strategy to train BianCang, which first injects domain-specific knowledge and then aligns it through targeted stimulation. Experimental results robustly demonstrate the effectiveness and superiority of BianCang.
Acknowledgment
@@
This work was supported by the National Natural Science Foundation of China (No.62376130 and No.62402258), Taishan Scholars Program (No.TSPD20240814 and No.TSQN202507242), Shandong Provincial Natural Science Foundation (No.ZR2022MF243), Program of New Twenty Policies for Universities of Jinan (No.202333008), Pilot Project for Integrated Innovation of Science, Education, and Industry of Qilu University of Technology (Shandong Academy of Sciences) (No.2025ZDZX01).
Clinical Relevance and Impact
The clinical relevance and impact of BianCang can be discussed from several key perspectives. First, BianCang significantly improves the accuracy of traditional Chinese medicine (TCM) syndrome differentiation and disease diagnosis in real-world scenarios, demonstrating its potential as a clinical diagnostic aid. When combined with retrieval-augmented generation (RAG) technology, BianCang provides TCM practitioners with reliable and accurate prescriptions, thus reducing the cognitive load associated with memorization.
Secondly, BianCang offers potential for enhancing the training and educational resources for TCM students and practitioners. With its superior performance in tasks involving real clinical scenarios and examinations, the model provides a reliable and accessible reference tool that can enrich the learning process.
Moreover, BianCang’s effectiveness in handling both pure TCM tasks and integrative medicine tasks (combining TCM with Western medicine) expands its applicability, demonstrating significant utility in contemporary healthcare environments where integrative approaches are increasingly valued.
Lastly, the open-source nature of BianCang ensures broad accessibility, enabling widespread clinical use, further validation, and continuous improvement by the global community. This democratization of specialized TCM knowledge potentially accelerates research, fosters collaboration, and elevates the overall standard of patient care in TCM practice.
Ethical Considerations
In the development and training of BianCang, rigorous measures were implemented to ensure the ethical use of patient data. All electronic medical records (EMRs) utilized in this study were derived from real-world patient data; however, comprehensive data anonymization and desensitization techniques were applied to safeguard patient privacy. Specifically, all personally identifiable information (PII), including names, addresses, contact details, and other sensitive data, was either removed or obfuscated. The training dataset was strictly anonymized, ensuring no information could be used to identify individual patients. This approach guarantees full compliance with ethical guidelines regarding patient privacy and confidentiality.
Additionally, the resources and data associated with this project are solely intended for academic research purposes. BianCang is designed as a language model-based intelligent assistant aimed at supporting research and providing insights into Traditional Chinese Medicine (TCM)-related topics. However, it is crucial to acknowledge the model’s inherent limitations, as it cannot guarantee the accuracy of all responses. BianCang is not a substitute for professional medical diagnosis or advice from licensed TCM or Western medicine practitioners. Users are strongly advised against relying exclusively on the model’s outputs for medical decision-making.
Given the potential serious consequences of inaccuracies in medical data, we emphasize the importance of exercising caution when interpreting and utilizing the information generated by the model. Users should always consult qualified healthcare professionals or seek medical attention at a hospital when necessary. BianCang is intended to complement, rather than replace, the expertise of medical professionals.
Furthermore, considering the potential safety risks in the application of TCM, we emphasize the following precautions: (1) Herbal compatibility safety, such as traditional contraindications including Shiba Fan (the ”Eighteen Incompatibilities”) and Shijiu Wei (the ”Nineteen Antagonisms”), should be carefully considered; (2) Medication safety during pregnancy, as certain herbs may pose risks to both the mother and fetus; (3) Diet-drug interactions, which are important in TCM practice and require attention to dietary restrictions and contraindications; and (4) Professional supervision, where all medication use should be guided by qualified TCM or medical professionals. Outputs from the model are intended for research purposes only and must not be used as a basis for actual medical treatment, to avoid potential serious consequences.